OpenStudy (thanatos2154):

derivative of ((x^2)+4)^x

OpenStudy (fibonaccichick666):

that's gonna be gross, ok, are you familiar with the chain rule?

OpenStudy (thanatos2154):

Yes.

OpenStudy (fibonaccichick666):

so, if you let \[x^2+4=u\] what is the derivative of that part?

OpenStudy (fibonaccichick666):

what is du?

OpenStudy (fibonaccichick666):

u'

OpenStudy (thanatos2154):

The derivative of u^x would be u^x right? No chain rule necessary.

OpenStudy (fibonaccichick666):

not yet

OpenStudy (fibonaccichick666):

first, we need du

OpenStudy (fibonaccichick666):

since, we have an exponent on a function

OpenStudy (thanatos2154):

I got as far as ln((x^2)+4)*(((x^2)+4)^x)

OpenStudy (thanatos2154):

But i looked it up on wolfram alpha, and it had some extra.

OpenStudy (fibonaccichick666):

That is not the derivative of the u I have defined.

OpenStudy (fibonaccichick666):

we will need it eventually, and it is easier to just not have to write all of that out so a substitution makes our lives easier

OpenStudy (thanatos2154):

Oh, ic. u' is 2x. right?

OpenStudy (fibonaccichick666):

good, ok so now we can say, we have \(u^x=f(x)\), you are correct that we will want to work with logs here to remove the exponent. So, which log is convenient to take in order to move the x down?

OpenStudy (thanatos2154):

ic, i normally use ln, my teacher says it doesn't matter.

OpenStudy (fibonaccichick666):

yep, it's convenient to work with so I prefer it as well, so now, remember you must take the ln of both sides, then differentiate, can you please show your steps here. (I can check for any misunderstandings this way)

OpenStudy (fibonaccichick666):

you can go step by step posting so I can nip any issues in the bud

OpenStudy (thanatos2154):